A farmer has 42 feet of fence with which to make a corral. If he arranges it into a rectangle that is twice as long as it is wide, what are the dimensions?
Found 2 solutions by savvyhush23, mykailao:
Answer by savvyhush23(50)

A farmer has 42 feet of fence with which to make a corral. If he arranges it into a rectangle that is twice as long as it is wide, what are the dimensions?
Perimeter = 42 ft.
twice as long as it is wide: L=2W
P = 2(L + W) = 2(2W + W) = 6W
42 = 6W
W = 7 ft. and L = 14 ft.

Answer by mykailao(6)

The farmer has 42 ft. of fence. In the problem, he arranges it into a rectangle.
P (perimeter) = 42ft.
L (length) = 2x
W (width) = x
Recall that the perimeter of a rectangle is P = 2 (L+W)
Subsitute now the P, L and W.
The new equation is 42 = 2 (2x + x)
Use the distributive property. Therefore it will become 42 = 4x + 2x
Find now the x.
4x + 2x = 42
6x = 42
x = 7
Answer:
L = 2(7) = 14ft.
W = x = 7ft.
Checking:
P = 2 (14 + 7)
= 2 (21)
= 42ft.

--There you go! :D